# FuseGPT: Learnable Layers Fusion of Generative Pre-trained Transformers

### v1

@@ Abstract
Generative Pre-trained Transformers (GPTs) have demonstrated remarkable performance across diverse domains through the extensive scaling of model parameters. Recent works observe the redundancy across the transformer blocks and develop compression methods by structured pruning of the unimportant blocks. However, such straightforward elimination will always provide irreversible performance degradation. In this paper, we propose FuseGPT, a novel methodology to recycle the pruned transformer blocks to further recover the model performance. Firstly we introduce a new importance detection metric, Macro Influence (MI), to detect the long-term influence of each transformer block by calculating their loss of information after removal. Then we propose group-level layers fusion, which adopts the parameters in layers of the unimportant blocks and injects them into the corresponding layers inside the neighboring blocks. The fusion is not one-off but through iterative parameter updates by lightweight group-level fine-tuning. Specifically, these injected parameters are frozen but weighted with learnable rank decomposition matrices to reduce the overhead during fine-tuning. Our approach not only works well on large language models but also on large multimodal models. The experiments have shown that, by using modest amounts of data, FuseGPT can outperform previous works in both perplexity and zero-shot task performance.
I Introduction
@@
Generative Pre-trained Transformers (GPTs) have proven to handle complex tasks and exhibit emergent abilities in various domains and tasks, especially when scaling up to billions of parameters [1, 2, 3, 4]. While GPTs have achieved unprecedented success, the growing complexity and size have brought increasing pressure on deployment in real-world scenarios, particularly in resource-constrained environments.
@@
To mitigate the challenges of the hardware requirements for GPTs deployment, model compression techniques are developed to produce compact models while preserving high performance. The techniques of model compression primarily fall into two categories: model pruning and quantization [5, 6, 7, 8, 9]. In this paper, we focus on model pruning, the technique to reduce the model size by eliminating redundant parameters. Pruning is mainly divided into two types: unstructured and structured [10, 11]. Unstructured pruning tries to remove individual weights, which can achieve higher performance but hinder the model acceleration due to the hardware-unfriendly sparse weights. Structured pruning, on the other hand, pruning the entire pre-defined model structure at once, which may slightly decrease accuracy but is more efficiently processed by hardware.
Recently, researchers have found that there exists redundancy across transformer blocks of the GPTs, i.e. some blocks seem to offer less contribution to the final results [12, 13, 14]. Some of them detect the redundancy by assessing the similarities between hidden states, and some others directly measure the changed distance to the hard label. After collecting the redundant blocks, they structured pruning the most unimportant of them from the model to prevent performance degradation as much as possible. However, directly discarding the blocks always leads to irreversible performance degradation. Traditional post-pruning fine-tuning may be helpful to recover the performance, but these techniques always need extremely large datasets and training resources. Therefore, efficient performance recovery techniques are urgently desired.
@@
Figure 1: Overview of pruning methods. Unstructured Pruning, Channel-wise Pruning, Block Pruning and our proposed FuseGPT.
@@
In this paper, we propose FuseGPT, a novel and efficient methodology for structured pruning. We observe the fact that despite some blocks being redundant, they still carry important pre-trained knowledge. It is intuitive to let it participate in the process of performance recovery. Therefore, we implement a full-scale technique to recycle and fuse the redundant blocks into other blocks before they are pruned. Firstly, a novel metric to detect the unimportant blocks is proposed that can achieve overall long-term influence assessment. During importance detection, we also consider the factor that how difficult it is to fuse them back, which is determined by the information loss caused by the block removal. Then we take steps to spread the parameters of layers inside the detected blocks to neighboring blocks. It is achieved by multiplying them with a learnable coefficient and injecting the weights into corresponding layers. Knowledge learning is then employed to update the parameters, which implicitly transfers the knowledge stored in the detected blocks. We ensure efficiency by restricting the fusion and parameter update processes within a partial group of blocks, and we constrain the coefficient matrix by decomposing it into low-rank matrices. Fig. 1 compares the proposed approach with other pruning methods.
@@
  • FuseGPT represents a novel methodology for performance recovery after redundant transformer blocks pruning. This approach not only presents high effectiveness but also shows outstanding efficiency.
  • To measure the long-term influence of block removal, this paper proposes a novel metric, Macro Inference (MI), for importance detection across the transformer blocks.
  • This paper introduces a novel technique to recycle the redundant blocks and fuse their layers into other blocks. This approach gives promising potential directions for future research on model compression.
  • Our experiments have demonstrated that FuseGPT can achieve state-of-the-art performance with superior perplexity on generation. When evaluating on zero-shot tasks, FuseGPT shows remarkable accuracy on both language and multimodal tasks.
II Related Work
@@
Quantization, which includes post-training quantization and quantization-aware training, has emerged as a highly favored approach in the domain of model compression. Conventionally, models are stored as floating-point values, but quantization offers an alternative by converting them into integer or other discrete representations. LLM.int8() has been firstly employed vector-wise quantization for the majority of the weight matrix [24]. Regarding the emergent outliers, they introduced a novel mixed-precision decomposition scheme, isolating the outlier feature dimensions into a 16-bit matrix multiplication. GPTQ proposed a more accurate data-aware approach by means of an approximate large-scale solver to minimize layer-wise L2 errors [25]. SmoothQuant [26] not only quantized both the weights and activations but also offline transferred the quantization difficulty from activations to weights through a mathematically equivalent transformation.
Pruning techniques, including both unstructured and structured ones, are employed to identify and remove redundant or less significant parameters from models, thus leading to a sparser weight matrix. ShortGPT[12] has put forward a straightforward layer removal approach that is based on Block Influence determined by the similarity between a layer’s input and output. Along this line, SLEB[14] has proposed a fundamental unit for pruning transformer blocks, enabling effective improvement in the processing speed of LLMs. In contrast, SliceGPT[27] replaces each weight matrix with a smaller (dense) matrix, reducing the embedding dimension of the network. FoldGPT[28] combines block removal and block parameter sharing. This work comprises two parts. Firstly, block importance, based on learnable gating parameters, determines the redundant layers according to the given removal rate. Secondly, for the retained blocks, a specially designed group parameter-sharing strategy is proposed to compress the number of parameters and slightly lower latency overhead.
Knowledge Distillation is widely used to transfer knowledge from a large model (teacher) to a smaller one (student) for improved efficiency, especially in the context of LLMs. DistilBERT [29] reduces the transformer’s layers in the teacher network by half and initializes the student by choosing one layer out of every two from the teacher. In contrast, MiniLM [30] simplifies the process by distilling knowledge solely from the self-attention module of the last Transformer block, thus alleviating the challenge of layer mapping. However, block removal and group parameter sharing based on the pre-trained model lead to additional performance degradation.
III Methodology
@@ III-A Preliminaries
Typically, researchers solve this by directly removing blocks in the GPT model to obtain a subset of transformer blocks $\mathcal{B}^{\prime}_{\mathcal{M}}\subseteq\mathcal{B}_{\mathcal{M}}$. Therefore, the problem becomes how to maintain the performance of the pruned model after blocks removal, and the key point is to remove the blocks that are relatively less important. We denote the process of choosing the unimportant block as importance detection. Since the residual path is added inside each transformation block, it is straightforward to detect the importance of blocks by measuring their similarity between the outputs of different transformer blocks during inference. Denote $\mathbf{X}_{i-1}$ and $\mathbf{X}_{i}$ ($i>=1$, and $\mathbf{X}_{0}$ denotes the input of $\mathcal{B}$) as the hidden states of two neighboring transformer blocks, ShortGPT [12] uses the metric Block Influence (BI) to measure the expectation of cosine similarity on the distribution for each token in a sequence, i.e. each $t$-th row of them, and the importance on $i$-th block is calculated as follows:
@@
A lower BI score indicates a small change by passing through the corresponding transformer block, which may represent lower importance during inference.
However, such an approach fails to consider the subsequent influence of the minor changes of the block removal can also be large. Therefore, SLEB [14] considers the overall model inference to evaluate the importance of the transformer block, i.e. calculating the loss of token prediction on the pruned model. Given a calibration tokenized sequence as $x=\{x_{1},...,x_{K}\}$ and denote the algorithm as SLEB score, it is calculated as follows:
@@
where the block to remove is iteratively selected by assessing the pruned model $\mathcal{M}^{\prime}_{i}$, i.e. previous pruned model $\mathcal{M}^{\prime}$ after removing the $i$-th block.
@@
Problem 1 (Importance Detection).
Given a generative pre-trained transformer model, the objective of importance detection is to develop a metric that identifies redundant or unimportant transformer blocks within the model, such that their removal minimizes performance degradation.
Despite a well-designed importance detection algorithm, the performance degradation of the pruned model is irreversible and becomes more severe as more blocks are removed. Therefore, methods to recover performance are necessary, a classical procedure in model compression, such as quantization-aware training and post-pruning fine-tuning. Besides direct re-training or fine-tuning of pruned models in structural pruning of transformer blocks, researchers have also developed specialized recovery methodologies. For example, in ShortGPT [12] and LLM-Streamline [31], they replace the removed block with a lightweight Multi-Layer Perceptron (MLP) that has much fewer parameters. Then the modified model is re-trained with mean squared error (MSE) loss.
@@
Problem 2 (Performance Recovery).
Given a pruned generative pre-trained transformer model after blocks removal, the objective of performance recovery is to develop an algorithm that can maximize the performance of the pruned model.
@@
III-B Importance detection via Macro Inference
As described in the preceding section, the importance of transformer blocks should be detected through overall model inference, to evaluate the long-term influence but not the local changes. Therefore, we propose a new metric, Macro Influence (MI), to measure the final influence of each block after removal. Denote the origin model and the pruned model with the removal of $i$-th block as $\mathcal{M}$ and $\mathcal{M}_{i}$, respectively. For convenience, we denote their last hidden states, i.e. the output of the last transformer block, as $\mathbf{X}_{\mathcal{M}}$ and $\mathbf{X}_{\mathcal{M}_{i}}$ respectively. By taking the cosine similarity between each row $t$ of them into consideration, the MI score of $i$-th block can then be calculated as follows:
@@ III-B Importance detection via Macro Inference
When removing one block to have little influence to the last hidden states, we expect that they will have high cosine similarity compared with the origin one and thus get a lower MI score.
The same is done for overall model inference, MI score works differently from the SELB score. Compared with directly using the token prediction results that calculates the loss on a hard label to measure the distance to the ground truth, we prefer considering the perturbation on the original results. By doing this, we can observe how one block matters to contribute to the original results and hence provide better instruction for performance recovery in later stages. As described by Hinton et al. [32], the soft targets have high entropy and hence can provide much more information than hard targets, thus by calculating the MI score we actually get the information loss of block removal. Moreover, it is expected that, when lower information loss is achieved, the difficulty of recovering the performance will also be decreased, thus beneficial for later performance recovery.
III-C Performance Recovery via Layers Fusion
Given the MI scores on each transformer block in $\mathcal{B}_{\mathcal{M}}$, we rank them in descending order to form a new set $\mathcal{B}^{MI}_{\mathcal{M}}=\{B^{MI}_{1},B^{MI}_{2},...,B^{MI}_{n}\}$. Firstly, we determine the removed block to be the the first-ranking block $B^{MI}_{1}$, i.e. the block with the highest MI score thus the most unimportant block.
Different from previous works, our goal is not simply abandoning it. Instead, we expect that it can contribute to the following performance recovery. According to this objective, we develop the idea to fuse the removed block to neighboring blocks, by doing this we hope that the functionalities belonging to it will be inherited by other blocks thus reducing the performance degradation.
Figure 2: In FuseGPT, we employ partial group to update the parameters.
@@ III-C Performance Recovery via Layers Fusion
$\text{where}\ \mathcal{I}$ =$\!=\!\begin{cases}\{p-\left\lfloor\frac{G}{2}\right\rfloor,...,p+\lceil\frac{G}{2}\rceil\}&\lceil\frac{G}{2}\rceil<p\leq n-\lceil\frac{G}{2}\rceil\\ \{1,...,G+1\}&1\leq p\leq\lceil\frac{G}{2}\rceil\\ \{n-G,...,n\}&n-\lceil\frac{G}{2}\rceil<p\leq n\end{cases}$
where we can observe that $|B_{partial}|=G+1$, i.e. $G$ neighboring blocks together with $B_{p}$, and we averagely split the $G$ blocks on both sides of $B_{p}$ in default. The strategy of arranging a partial group to do fusion and fine-tuning offers dual advantages. On the one hand, it maintains computational efficiency for contemporary large-scale GPTs. On the other hand, the blocks adjacent to the pruned block are expected to have similar functionalities, which meets our objective since fusing similar blocks will reduce the difficulty.
In each transformer block, we treat the weights on linear layers as the fundamental unit for fusion, which account for most of the parameters and take primary responsibility for functionality. Suppose we want to fuse the block to prune $B_{p}$ into one of the blocks $B_{i}$ inside the partial group $\mathcal{B}_{partial}$, we denote a linear layer in $B_{p}$ as $l_{p,j}$ and the corresponding layer $l_{i,j}$ in $B_{i}$ that serves the same functional role, e.g. both are the first linear layer of the feed-forward module. Denote $\mathbf{W}_{i,j}\in\mathbb{R}^{d\times k}$ and $\mathbf{W}_{p,j}\in\mathbb{R}^{d\times k}$ as the weights of $l_{i}$ and $l_{p}$, and we employ a learnable matrix $\mathbf{C}\in\mathbb{R}^{d\times k}$ as coefficient for $\mathbf{W}_{p,j}$. Then they are weighted added together as the weight of the fused layer $l_{i,j}^{fused}$:
@@
We initialize $\mathbf{C}_{right}$ with Kaiming initialization [34] and zero for $\mathbf{C}_{left}$ to build a good start-point of learning. During fine-tuning, the $\mathbf{W}_{p,j}$ is frozen and the gradient updates is conducted on $\mathbf{C}_{left}$, $\mathbf{C}_{right}$ and $\mathbf{W}_{i,j}$. We keep doing the fusion for all the linear layers in the Block $B_{i}$ to obtain $B_{i}^{fused}$, and then doing fusion for all the blocks inside $\mathcal{B}_{partial}$ (except $B_{p}$). Finally we remove $B_{p}$ from the group and obtain the fused partial group $\mathcal{B}_{partial}^{fused}=\{B_{1}^{fused},...,B_{G}^{fused}\}$ and $|\mathcal{B}_{partial}^{fused}|=G$.
@@
3:    $\mathcal{B}_{\mathcal{M}}\longleftarrow\{B_{1},...,B_{n-i}\}$
@@
6:         $S\longleftarrow\text{MI}_{j}(\mathcal{M},\mathcal{C})$triangleright Equation 3
@@
13:    $\mathcal{M}\leftarrow layers\_fusion(\mathcal{M},B_{p})$triangleright Algorithm 2
@@
2:$\mathcal{B}_{partial}\leftarrow get\_parital\_group(\mathcal{M},B_{p},G)$triangleright Equation 4
@@
5:         $l_{p,j}\longleftarrow$ layer to fuse in $B_{p}$
6:         $\mathbf{W}_{i,j},\mathbf{W}_{p,j}\longleftarrow$ weights of $l_{i,j},l_{p,j}$
7:         $\mathbf{C}=\mathbf{C}_{left}\mathbf{C}_{right}\longleftarrow$ low-rank coefficient
8:         $\mathbf{W}^{fused}_{i,j}\longleftarrow\mathbf{W}_{i,j}+\mathbf{C}\odot\mathbf{W}_{p,j}$triangleright Equation 5
9:         $l_{i,j}^{fused}\longleftarrow$ fused layer with weight $\mathbf{W}^{fused}_{i,j}$
@@
11:    $\mathcal{B}_{i}^{fused}\longleftarrow$ fused block
@@
13:$\mathcal{B}_{partial}^{fused}\longleftarrow\{B_{1}^{fused},...,B_{G}^{fused}\}$
14:Compute the KL divergence loss $\mathcal{L_{KL}}$ with $\mathcal{D}$triangleright Equation 9
15:Update $\mathcal{B}_{partial}^{fused}$ by minimizing $\mathcal{L_{KL}}$
16:$\mathcal{M}\longleftarrow group\_replace(\mathcal{M},\mathcal{B}_{partial},\mathcal{B}_{partial}^{fused})$
17:fused model $\mathcal{M}$
It should be noted that the process from importance detection to performance detection is conducted iteratively, i.e. the blocks are fused one by one until the predefined pruning rate is achieved. The reason is that the current importance detection result is based on the current state of overall model inference, but once a block is removed the states will also change, which will also change the rank of block importance. Such influence will even be more significant after we do the fusion and knowledge-learning processes.
We summarize the overall algorithm of FuseGPT in Algorithm 1. In each iteration, we calculate the MI score to detect the most unimportant block. Then we apply group-level layers fusion as in Algorithm 2. We create a partial group of blocks around the detected block to prune. Then reparameterization is done on the layers of the partial group by fusing them with the corresponding layers in the detected block. In the end, lightweight partial group fine-tuning is performed to learn the fusion from knowledge loss.
IV Experiments
In this section, we first introduce the experimental setting for FuseGPT implementation. Then we present our main results, from perplexity to zero-shot task results after model pruning. We also conduct ablation studies to highlight our methodologies.
TABLE I: Perplexity results on WikiText-2 and C4 datasets. Randomly select samples from WikiText-2 training dataset as calibration data. We omit the results of SliceGPT on LLaVA models due to their current lack of implementation support.
@@ IV Experiments
TABLE II: Zero-shot task results for language models. Randomly select samples from WikiText-2 training dataset as calibration data.

### v2

@@ Abstract
Generative Pre-trained Transformers (GPTs) have demonstrated remarkable performance across diverse domains, largely due to the extensive scaling of model parameters. Recent works have observed redundancy within transformer blocks and developed compression methods by structured pruning of less important blocks. However, such direct removal often leads to irreversible performance degradation. In this paper, we propose FuseGPT, a novel methodology designed to recycle pruned transformer blocks, thereby recovering the model’s performance. Firstly, we introduce a new importance detection metric, Macro Influence (MI), which evaluates the long-term impact of each transformer block by quantifying the information loss incurred upon its removal. Next, we propose group-level layer fusion, which leverages the parameters from layers of less important blocks and integrates them into the corresponding layers of neighboring blocks. This fusion process is not a one-time operation but is refined through iterative parameter updates by lightweight group-level fine-tuning. Specifically, the injected parameters are frozen but are weighted with learnable rank decomposition matrices to reduce the computational overhead during fine-tuning. Our approach not only works well for large language models but also for large multimodal models. Experimental results indicate that, even with modest amounts of data, FuseGPT surpasses previous methods in both perplexity and zero-shot task performance.
1 Introduction
@@
Generative Pre-trained Transformers (GPTs) have demonstrated remarkable capabilities in handling complex tasks and exhibiting emergent abilities in various domains, especially when scaled to billions of parameters [3, 53, 43, 26]. Despite their unprecedented success, the increasing complexity and size of GPTs have introduced significant challenges for deployment in real-world scenarios, particularly in resource-constrained environments.
@@
To address the hardware demands associated with deploying GPTs, model compression techniques are developed to produce more compact models while preserving high performance. These techniques primarily fall into two categories: model pruning and quantization [22, 12, 15, 28, 11]. This paper focuses on model pruning, a technique aimed at reducing model size by eliminating redundant parameters. Pruning is mainly categorized into two types: unstructured pruning and structured pruning [7, 45]. Unstructured pruning targets at removing individual weights, which can achieve higher performance but often results in hardware-unfriendly sparse weights, limiting acceleration potential. Structured pruning, in contrast, removes entire pre-defined structures (e.g., layers or blocks) at once, which may lead to a slight reduction in accuracy but is more hardware-efficient.
Recent studies have revealed that redundancy exists across transformer blocks in GPTs, meaning that certain blocks contribute less significantly to the final outcomes [34, 19, 41]. Some existing methods detect this redundancy by analyzing the similarities between hidden states, while others directly measure the changes in distance to the hard labels. Once redundant blocks are identified, structured pruning is applied to remove the least important ones, aiming to minimize the performance degradation. However, simply discarding these blocks often results in irreversible performance loss. While traditional post-pruning fine-tuning can help recover the performance, they typically require extensive datasets and ubstantial computational resources. As a result, there is a pressing need for more efficient methods to restore model performance without such heavy demands.
@@
Figure 1: Overview of pruning methods. Unstructured Pruning, Channel-wise Pruning, Block Pruning, and our proposed FuseGPT.
@@
In this paper, we propose FuseGPT, a novel and efficient methodology for structured pruning. We observe that even though some transformer blocks may be redundant, they still carry valuable pre-trained knowledge. Rather than directly discarding these blocks, we propose to recycle and fuse their knowledge into other blocks before their removal, enabling a more effective recovery of model performance. Firstly, we introduce a novel metric designed to identify less important blocks by assessing their long-term influence on the model’s performance. During importance detection, we also evaluate the difficulty of fusing each block back into the model, which is determined by the extent of information loss caused by its removal. Next, we take steps to spread the parameters from the layers of the detected redundant blocks to neighboring blocks. This is achieved by multiplying the redundant block’s parameters with a learnable coefficient and injecting the resulting weights into the corresponding layers of neighboring blocks. This process can be viewed as a knowledge transfer mechanism, where the knowledge stored in the redundant blocks is implicitly transferred to their neighboring blocks. To maintain efficiency, we restrict the fusion and parameter update processes to a partial group of blocks rather than the entire model. Additionally, we optimize the learnable coefficient matrix by decomposing it into low-rank matrices to reduce computational overhead. Figure 1 compares the proposed approach with other pruning methods.
@@
  • FuseGPT offers a novel methodology for performance recovery following the pruning of redundant transformer blocks. It is characterized by a new metric, Macro Inference (MI), to measure the long-term impact of block removal and identify the importance of each transformer block to minimize redundancy.
  • The methodology also introduces a novel technique for recycling redundant blocks by fusing their layers into neighboring blocks, which opens new avenues for future research in model compression.
Our experimental results unequivocally demonstrate the exceptional effectiveness and efficiency of FuseGPT. It achieves state-of-the-art performance with superior perplexity in generation tasks. Furthermore, in zero-shot task evaluations, FuseGPT exhibits remarkable accuracy across both language-based and multimodal tasks.
2 Related Work
@@
Pruning techniques. Pruning, including both unstructured and structured ones, are employed to identify and remove redundant or less significant parameters from models, thus leading to a sparser weight matrix. ShortGPT [34] has put forward a straightforward layer removal approach that is based on Block Influence determined by the similarity between a layer’s input and output. Along this line, SLEB [41] has proposed a fundamental unit for pruning transformer blocks, enabling effective improvement in the processing speed of LLMs. In contrast, SliceGPT [1] replaces each weight matrix with a smaller (dense) matrix, reducing the embedding dimension of the network. FoldGPT [27] combines block removal and block parameter sharing. This work comprises two parts. Firstly, block importance, based on learnable gating parameters, determines the redundant layers according to the given removal rate. Secondly, for the retained blocks, a specially designed group parameter-sharing strategy is proposed to compress the number of parameters and slightly lower latency overhead. LaCo [48] uses layer merging to compress the model.
Knowledge Distillation. Knowledge distillation is widely used to transfer knowledge from a large model (teacher) to a smaller one (student) for improved efficiency, especially in the context of LLMs. DistilBERT [40] reduces the transformer’s layers in the teacher network by half and initializes the student by choosing one layer out of every two from the teacher. In contrast, MiniLM [44] simplifies the process by distilling knowledge solely from the self-attention module of the last Transformer block, thus alleviating the challenge of layer mapping. However, block removal and group parameter sharing based on the pre-trained model lead to additional performance degradation.
3 Methodology
@@ 3.1 Preliminaries
Typically, researchers solve this by directly removing blocks in the GPT model to obtain a subset of transformer blocks $\mathcal{B}^{\prime}_{\mathcal{M}}\subseteq\mathcal{B}_{\mathcal{M}}$. Therefore, the problem becomes how to maintain the performance of the pruned model after blocks removal, and the key point is to remove the blocks that are relatively less important. We denote the process of choosing the unimportant block as importance detection. Since the residual path is added inside each transformation block, it is straightforward to detect the importance of blocks by measuring their similarity between the outputs of different transformer blocks during inference. Denote $\mathbf{X}_{i-1}$ and $\mathbf{X}_{i}$ ($i>=1$, and $\mathbf{X}_{0}$ denotes the input of $B_{1}$) as the hidden states of two neighboring transformer blocks, ShortGPT [34] uses the metric Block Influence (BI) to measure the expectation of cosine similarity on the distribution for each token in a sequence, i.e. each $t$-th row of them, and the importance on $i$-th block is calculated as follows:
@@
A lower BI score indicates a smaller change by passing through the corresponding transformer block, which implies lower importance during inference.
However, such an approach fails to consider that the subsequent influence of the minor changes of the block removal can also be large. Therefore, SLEB [41] considers the overall model inference to evaluate the importance of the transformer block, i.e., calculating the loss of token prediction on the pruned model. Given a calibration tokenized sequence as $x=\{x_{1},...,x_{K}\}$ and denote the algorithm as SLEB score, it is calculated as follows:
@@
where the block to remove is iteratively selected by assessing the pruned model $\mathcal{M}^{\prime}_{i}$, i.e., the previous pruned model $\mathcal{M}^{\prime}$ after removing the $i$-th block.
@@
Problem 1 (Importance Detection)
Given a GPT model, the objective of importance detection is to develop a metric that identifies redundant or unimportant transformer blocks within the model, such that their removal minimizes performance degradation.
Despite the use of well-designed importance detection algorithms, the performance degradation of pruned models remains irreversible and worsens as more blocks are removed. Therefore, performance recovery methods are essential, a classical procedure in model compression, such as quantization-aware training and post-pruning fine-tuning. Besides direct re-training or fine-tuning of pruned models in structural pruning of transformer blocks, researchers have also developed specialized recovery methodologies. For example, in ShortGPT [34] and LLM-Streamline [4], they replace the removed block with a lightweight Multi-Layer Perceptron (MLP) that has much fewer parameters. Then, the modified model is re-trained with mean squared error (MSE) loss.
@@
Problem 2 (Performance Recovery)
Given a pruned GPT model after block removal, the objective of performance recovery is to develop an algorithm that can maximize the performance of the pruned model.
@@
3.2 Importance detection via Macro Inference
As discussed in the previous section, the importance of transformer blocks should be assessed through overall model inference to evaluate their long-term influence rather than localized changes. To this end, we introduce a new metric, Macro Influence (MI), designed to measure the final impact of each block after its removal. Denote the original model and the pruned model with the $i$-th block removed by $\mathcal{M}$ and $\mathcal{M}_{i}$, respectively. For simplicity, we denote their last hidden states, i.e., the output of the final transformer block, by $\mathbf{X}_{\mathcal{M}}$ and $\mathbf{X}_{\mathcal{M}_{i}}$, respectively. By considering the cosine similarity between each row $t$ of these hidden states, the MI score of the $i$-th block is calculated as follows:
@@ 3.2 Importance detection via Macro Inference
When the removal of a block has little impact on the last hidden states, the cosine similarity for each row will be high, resulting in a low MI score.
Unlike the SLEB score, which relies on token prediction results and measures loss based on hard labels to determine the distance to the ground truth, the MI score focuses on the perturbation of the original results. This approach allows us to better understand how each block contributes to the model’s output, providing more precise guidance for performance recovery in subsequent stages. As emphasized by Hinton et al. [14], the soft targets have high entropy and thus offer significantly more information than hard targets. By calculating the MI score, we quantify the information loss caused by block removal. Importantly, achieving lower information loss during this process is expected to reduce the difficulty of performance recovery, making it more efficient and effective. This makes the MI score not only a robust metric for importance detection but also a valuable tool for optimizing performance recovery strategies.
3.3 Performance Recovery via Layers Fusion
Given the MI scores on each transformer block in $\mathcal{B}_{\mathcal{M}}$, we rank them in ascending order to form a new set $\mathcal{B}^{MI}_{\mathcal{M}}=\{B^{MI}_{1},B^{MI}_{2},...,B^{MI}_{n}\}$. he block with the lowest MI score, denoted as $B^{MI}_{1}$, is identified as the least important block and is selected for removal.
Unlike previous approaches that simply discard such blocks, our goal is to recycle and reuse their knowledge to aid in subsequent performance recovery. To achieve this, we propose fusing the removed block into its neighboring blocks. This fusion process ensures that the functionalities of the removed block are inherited by other blocks, thereby mitigating performance degradation. By integrating the knowledge from the redundant block into its neighbors, we aim to preserve its contributions to the model while reducing redundancy and maintaining efficiency.
Figure 2: In FuseGPT, we employ a partial group to update the parameters.
@@ 3.3 Performance Recovery via Layers Fusion
$\text{where}\ \mathcal{I}$ =$\!=\!\begin{cases}\{p-\left\lfloor\frac{G}{2}\right\rfloor,...,p+\lceil\frac{G}{2}\rceil\},&\lceil\frac{G}{2}\rceil<p\leq n-\lceil\frac{G}{2}\rceil,\\ \{1,...,G+1\},&1\leq p\leq\lceil\frac{G}{2}\rceil,\\ \{n-G,...,n\},&n-\lceil\frac{G}{2}\rceil<p\leq n.\end{cases}$
Here, we can observe that $|B_{partial}|=G+1$, i.e., $G$ neighboring blocks together with $B_{p}$, and we averagely split the $G$ blocks on both sides of $B_{p}$ in default. The strategy of arranging a partial group to do fusion and fine-tuning offers dual advantages. On the one hand, it maintains computational efficiency for contemporary large-scale GPTs. On the other hand, the blocks adjacent to the pruned block are expected to have similar functionalities, which meets our objective since fusing similar blocks will reduce the difficulty.
In each transformer block, we treat the weights on linear layers as the fundamental unit for fusion, which accounts for most of the parameters and takes primary responsibility for functionality. Suppose we want to fuse the block to prune $B_{p}$ into one of the blocks $B_{i}$ inside the partial group $\mathcal{B}_{partial}$, we denote a linear layer in $B_{p}$ as $l_{p,j}$ and the corresponding layer $l_{i,j}$ in $B_{i}$ that serves the same functional role, e.g. both are the first linear layer of the feed-forward module. Denote $\mathbf{W}_{i,j}\in\mathbb{R}^{d\times k}$ and $\mathbf{W}_{p,j}\in\mathbb{R}^{d\times k}$ as the weights of $l_{i}$ and $l_{p}$, and we employ a learnable matrix $\mathbf{C}\in\mathbb{R}^{d\times k}$ as coefficient for $\mathbf{W}_{p,j}$. Then, they are weighted and added together as the weight of the fused layer $l_{i,j}^{fused}$:
@@
We initialize $\mathbf{C}_{right}$ with Kaiming initialization [13] and zero for $\mathbf{C}_{left}$ to build a good starting point for learning. During fine-tuning, the $\mathbf{W}_{p,j}$ is frozen and the gradient updates is conducted on $\mathbf{C}_{left}$, $\mathbf{C}_{right}$ and $\mathbf{W}_{i,j}$. We keep doing the fusion for all the linear layers in the Block $B_{i}$ to obtain $B_{i}^{fused}$, and then doing fusion for all the blocks inside $\mathcal{B}_{partial}$ (except $B_{p}$). Finally we remove $B_{p}$ from the group and obtain the fused partial group $\mathcal{B}_{partial}^{fused}=\{B_{1}^{fused},...,B_{G}^{fused}\}$ and $|\mathcal{B}_{partial}^{fused}|=G$.
@@
3:    $\mathcal{B}_{\mathcal{M}}\longleftarrow\{B_{1},...,B_{n-i}\}$;
@@
6:         $S\longleftarrow\text{MI}_{j}(\mathcal{M},\mathcal{C})$; $\triangleright$ Eq. (3)
@@
13:    $\mathcal{M}\leftarrow\texttt{layers\_fusion}(\mathcal{M},B_{p})$; $\triangleright$ Eq. (2)
@@
2:$\mathcal{B}_{partial}\leftarrow\texttt{get\_parital\_group}(\mathcal{M},B_{p},G)$triangleright Eq. (4);
@@
5:         $l_{p,j}\longleftarrow$ layer to fuse in $B_{p}$;
6:         $\mathbf{W}_{i,j},\mathbf{W}_{p,j}\longleftarrow$ weights of $l_{i,j},l_{p,j}$;
7:         $\mathbf{C}=\mathbf{C}_{left}\mathbf{C}_{right}\longleftarrow$ low-rank coefficient;
8:         $\mathbf{W}^{fused}_{i,j}\longleftarrow\mathbf{W}_{i,j}+\mathbf{C}\odot\mathbf{W}_{p,j}$; $\triangleright$ Eq. (5)
9:         $l_{i,j}^{fused}\longleftarrow$ fused layer with weight $\mathbf{W}^{fused}_{i,j}$;
@@
11:    $\mathcal{B}_{i}^{fused}\longleftarrow$ fused block;
@@
13:$\mathcal{B}_{partial}^{fused}\longleftarrow\{B_{1}^{fused},...,B_{G}^{fused}\}$;
14:Compute the KL divergence loss $\mathcal{L_{KL}}$ with $\mathcal{D}$; $\triangleright$ Eq. (9)
15:Update $\mathcal{B}_{partial}^{fused}$ by minimizing $\mathcal{L_{KL}}$;
16:$\mathcal{M}\longleftarrow\texttt{group\_replace}(\mathcal{M},\mathcal{B}_{partial},\mathcal{B}_{partial}^{fused})$;
17:fused model $\mathcal{M}$;
It should be noted that the process from importance detection to performance detection is conducted iteratively, i.e. the blocks are fused one by one until the predefined pruning rate is achieved. The reason is that the current importance detection result is based on the current state of overall model inference, but once a block is removed, the states will also change, which will also change the rank of block importance. Such influence will even be more significant after we do the fusion and knowledge-learning processes.
We summarize the overall algorithm of FuseGPT in Algorithm 1. In each iteration, we calculate the MI score to detect the most unimportant block. Then we apply group-level layers fusion as in Algorithm 2. We create a partial group of blocks around the detected block to prune. Then, reparameterization is done on the layers of the partial group by fusing them with the corresponding layers in the detected block. In the end, lightweight partial group fine-tuning is performed to learn the fusion from knowledge loss.
Table 1: Perplexity results on WikiText-2 and C4 datasets. Randomly select samples from WikiText-2 training dataset as calibration data. We omit the results of SliceGPT on LLaVA models due to their current lack of implementation support.
@@
4 Experiments
Table 2: Zero-shot task results for language models. Randomly select samples from WikiText-2 training dataset as calibration data.